OpenStudy (anonymous):

Eighteen is 60% of what number?

OpenStudy (anonymous):

I think 25

OpenStudy (anonymous):

\[18 = \frac{60}{100}(x)\]

OpenStudy (anonymous):

`of` = \(\times\) `is` means =

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Just solve for x.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is it 30

OpenStudy (anonymous):

`%` means \(\frac{}{100}\)

OpenStudy (anonymous):

`what number` becomes a variable, like \(x\).

OpenStudy (anonymous):

The number that 18 is 60% of

OpenStudy (anonymous):

Yes. \(30\cdot 60\%=18\).

OpenStudy (anonymous):

so it is 30

OpenStudy (anonymous):

ok thank you